Maynard Steel plans to pay a dividend of $3.14 this year. The company has an expected earnings growth rate of 3.7% per year and an equity cost of capital of 10 796. a. Assuming Maynards dividend payout rate and expected growth rate remain constant, and that the firm does not issue or repurchase shares, estimate Maynards share price b. Suppose Maynard decides to pay a dividend of $1.07 this year and use the remaining $2.07 per share to repurchase shares. If Maynards total payout rate remains constant, estimate Maynards share price c. If Maynard maintains the dividend and total payout rate in (b), at what rate are Maynards dividends and eamings per share expected to
Answer #1

a. Maynard's share price=Dividend/(cost of capital-growth rate)

=3.14/(10.7-3.7)=$44.86

b. Using total payout model , as the rate remains constant.

Share price= 3.14/(10.7-3.7)=$44.86

c. Growth rate= Cost of capital-(Dividend/share price)

=10.7-(1.7/44.86)=8.31%

Please rate my answer